### v1

# Differentially Private Low-Rank Adaptation of Large Language Model Using Federated Learning

Abstract.
The surge in interest and application of large language models (LLMs) has sparked a drive to fine-tune these models to suit specific applications, such as finance and medical science. However, concerns regarding data privacy have emerged, especially when multiple stakeholders aim to collaboratively enhance LLMs using sensitive data. In this scenario, federated learning becomes a natural choice, allowing decentralized fine-tuning without exposing raw data to central servers. Motivated by this, we investigate how data privacy can be ensured in LLM fine-tuning through practical federated learning approaches, enabling secure contributions from multiple parties to enhance LLMs. Yet, challenges arise: 1) despite avoiding raw data exposure, there is a risk of inferring sensitive information from model outputs, and 2) federated learning for LLMs incurs notable communication overhead. To address these challenges, this article introduces DP-LoRA, a novel federated learning algorithm tailored for LLMs. DP-LoRA preserves data privacy by employing a Gaussian mechanism that adds noise in weight updates, maintaining individual data privacy while facilitating collaborative model training. Moreover, DP-LoRA optimizes communication efficiency via low-rank adaptation, minimizing the transmission of updated weights during distributed training. The experimental results across medical, financial, and general datasets using various LLMs demonstrate that DP-LoRA effectively ensures strict privacy constraints while minimizing communication overhead.
Federated learning, differential privacy, domain-specific LLM, LLM privacy, medical LLM, financial LLM, efficient fine-tuning
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
1. Introduction
Recent strides made in large language models (LLMs), like ChatGPT (Ouyang et al., 2022) and GPT-4 (OpenAI, 2023), have significantly propelled the field of natural language processing (NLP) forward. Their remarkable success has naturally sparked interest in domain-specific applications, such as finance (Wu et al., 2023; Liu et al., 2023a; Yang et al., 2023; Liu et al., 2023b; Xie et al., 2023), medical science (Singhal et al., 2023), law (Nguyen, 2023), biology (Luo et al., 2022) and beyond. However, employing general-purpose LLMs directly in these specialized domains may result in subpar performance due to the inherent discrepancy between general and domain-specific texts (Zha et al., 2023a, b). Consequently, there arises an increasing demand to fine-tune LLMs for specific domains (Hu et al., 2021; Dodge et al., 2020; Yu et al., 2021; Ding et al., 2023). Notably, OpenAI has released ChatGPT fine-tuning API [^1] https://platform.openai.com/docs/guides/fine-tuning, empowering users to tailor ChatGPT by fine-tuning it with their own datasets, effectively adapting it to their unique applications.
Despite the promise of LLM fine-tuning, there is a growing concern surrounding the potential leakage of private data while utilizing LLM (Li et al., 2023a; Sebastian, 2023), particularly within high-stakes domains. Consider, for instance, a scenario where multiple hospitals collaborate to enhance LLMs by fine-tuning them using their medical data. While this collaborative effort could significantly enhance model performance and offer substantial benefits to patients, hospitals may refrain from committing to this practice due to concerns about exposing sensitive patient information to other institutions and the public. In this scenario, federated learning strategies (Li et al., 2020a; Zhang et al., 2021) emerge as a natural choice, enabling distributed fine-tuning of the model with locally processed data, avoiding raw data exposure to servers. Therefore, we are motivated to study the following research question: How can we guarantee data privacy in LLM fine-tuning with a practical federated learning approach, enabling each stakeholder to securely contribute data and derive collective advantages from an improved LLM?
However, accomplishing the aforementioned objective presents non-trivial challenges. Firstly, despite federated learning ensuring no direct exposure to raw data, there remains a risk for attackers to potentially deduce training data from model outputs. This vulnerability arises because LLMs can accidentally disclose sensitive information through their responses (Carlini et al., 2021; Kim et al., 2023; Li et al., 2023a). For instance, reports indicate that malicious adversaries could exploit the New Bing to link victims’ personally identifiable information using partial information (Li et al., 2023b). Secondly, naively applying federated learning to fine-tune LLMs can trigger significant communication overhead. This is because model updates from diverse decentralized sources need iterative aggregation, and frequent transmission of these updates for a complex model like LLM is extremely expensive. Despite strategies to mitigate communication challenges in federated learning (Konečnỳ et al., 2016; Hamer et al., 2020; Lan et al., 2023), a notable trade-off persists: while these methods alleviate communication bottlenecks, they cannot ensure concrete guarantees in protecting the privacy of the training data.
To address the above challenges, in this paper, we introduce Differentially Private Low-Rank Adaptation (DP-LoRA), a communication-efficient federated learning algorithm designed for LLMs with privacy guarantees. Ensuring data privacy, DP-LoRA employs a Gaussian mechanism, introducing random noise in weight updates to ensure minimal changes in publicly visible information when an individual in the dataset changes. This approach prevents any single sample from significantly impacting the output, thereby thwarting attackers’ attempts to infer private information corresponding to specific individuals. We have substantiated our algorithm’s adherence to the $(\epsilon,\delta)$-differential privacy property, a widely accepted criterion in differential privacy. Addressing the communication overhead, DP-LoRA leverages the low-rank characteristic inherent in LLMs (Hu et al., 2021) by incorporating a smaller set of new weights into the model, focusing fine-tuning solely on these parameters. Consequently, only a limited number of updated weights necessitate transmission during distributed training, resulting in substantial reductions in communication costs. In summary, this article contributes in the following ways:
  • We identify the need for preserving privacy in LLM fine-tuning and frame it as a federated learning paradigm, seeking to empower each stakeholder to securely contribute data and collectively benefit from enhancing the LLM.
  • We present DP-LoRA, a novel federated learning algorithm specifically crafted for LLMs. DP-LoRA ensures differential privacy with privacy guarantees, while also reducing communication overhead via low-rank adaptation.
  • Through experiments across medical, financial, and general datasets using four LLMs, we demonstrate the efficacy of DP-LoRA. The results consistently indicate its ability to maintain strict privacy constraints while minimizing communication overhead.
The rest of the paper is organized as follows: Section 2 provides a review of related work in LLM privacy, domain-specific LLM, parameter-efficient tuning for LLMs, federated learning, and differential privacy. Section 3 delves into the foundational principles of differential privacy, elucidating its crucial role in preserving data privacy during training. Moving forward, Section 4 details the proposed DP-LoRA algorithm for LLMs, which integrates low-rank adaptation with differential privacy to address communication costs while maintaining data privacy. Subsequently, Section 5 presents the performance evaluation, analyzing and comparing the effectiveness of the proposed methods across general, medical, and finance datasets using various LLMs. Finally, Section 6 concludes this paper.
2. Related Work
2.1. Privacy Issues of LLMs
With the rise of LLMs, the issue of privacy has garnered significant attention, where it has been frequently shown that attackers can exploit model responses to infer sensitive information (Carlini et al., 2021; Kim et al., 2023; Li et al., 2023a, b; Carlini et al., 2019). This surge in concern has sparked a growing interest in preserving privacy within LLMs (Shi et al., 2021; Li et al., 2023a; Yu et al., 2021; Anil et al., 2021; Hoory et al., 2021; Li et al., 2021). However, current efforts tend to either focus solely on the pre-training phase of LLMs or introduce substantial communication overhead due to a large number of weight updates, significantly limiting the algorithm’s practical application. Furthermore, these efforts primarily concentrate on general text data rather than domain-specific datasets. To tackle these limitations, we propose an efficient privacy-preserving fine-tuning algorithm for LLMs, aiming to ensure privacy with minimal communication overhead. We also showcase its effectiveness across medical, financial, and general datasets.
2.2. From General-Purpose LLMs to Domain-specific LLMs
The emergence of LLMs marks a significant shift in NLP (Zhao et al., 2023). Examples such as GPT-2/3/4 (Radford et al., 2019; Brown et al., 2020; OpenAI, 2023), GPT-NeoX (Black et al., 2022), BERT (Devlin et al., 2018), PaLM (Chowdhery et al., 2022), BLOOM (Scao et al., 2022), OPT (Zhang et al., 2022), and LLaMA (Touvron et al., 2023) showcase exceptional language comprehension and generation abilities, excelling in various practical tasks. Recently, there has been a growing trend in crafting LLMs specialized for particular domains. In the following, we delve into recent LLMs efforts on medical and financial domains.
In the medical field, Med-PaLM (Singhal et al., 2023) emerges as a medical-specific LLM refined through instruction tuning using medical benchmarks, showcasing performance akin to clinicians. Additionally, BioGPT (Luo et al., 2022), a generative Transformer language model specifically pre-trained on expansive biomedical literature, demonstrates encouraging performance across diverse biomedical datasets. These initiatives across different domains have underscored the efficacy of domain-specific LLMs.
In the finance domain, the advent of FinLLMs starts with BloombergGPT (Wu et al., 2023), a model trained on a blend of financial and general datasets. It demonstrates promise across crucial financial tasks like forecasting, sentiment analysis, and risk assessment. However, the closed-source nature of BloombergGPT’s dataset poses obstacles to advancing FinLLMs, urging the exploration of cost-effective methods for domain adaptation. Recent endeavors, like FinGPT (Liu et al., 2023a), aim to democratize access to vast internet-scale data for training financial LLMs by developing diverse data pipelines to acquire and refine high-quality training data. Additionally, Pixiu (Xie et al., 2023) emerges as a tool for benchmarking LLMs in financial tasks, leveraging superior datasets to enhance the evaluation of model efficacy in this domain.
Despite the promise of domain-specific LLMs, the privacy issue remains prevalent. This concern becomes notably amplified in high-stakes domains such as finance and medical science. In this article, we aim to address this challenge by proposing efficient algorithms for preserving privacy during the fine-tuning process of domain-specific LLMs.
2.3. Parameter-Efficient Tuning of LLMs
The implementation of efficient fine-tuning methods for LLMs has garnered significant attention (Hu et al., 2021; Dettmers et al., 2023; Lester et al., 2021; Sung et al., 2022; Liu et al., 2023c; Zaken et al., 2021; Karimi Mahabadi et al., 2021). Notably, some widely recognized approaches include the utilization of adapters (Houlsby et al., 2019; Karimi Mahabadi et al., 2021), which embed a compact module within transformer blocks, allowing selective updates while keeping other parameters fixed. Similarly, the Low-Rank Adaptation (LoRA) technique (Hu et al., 2021) integrates trainable rank decomposition matrices into transformer blocks, facilitating updates solely to these matrices while preserving the remaining parameters. However, these parameter-efficient tuning methods do not consider data privacy. In this article, we propose leveraging federated learning and differential privacy to enable secure fine-tuning of LLMs.
2.4. Federated Learning and Differential Privacy
Federated Learning, a recently proposed training approach, focuses on safeguarding dataset privacy during distributed training. Early implementations involved employing model averaging to minimize communication overhead (McMahan et al., 2017). Building upon this foundation, recent advancements have introduced various algorithms, including variance reduction methods (Das et al., 2022) and penalty-based approaches (Li et al., 2020b), in both homogeneous (Woodworth et al., 2020) and heterogeneous (Karimireddy et al., 2020) settings. Notably, the integration of differential privacy into federated learning has gained traction, offering a theoretical framework to govern privacy boundaries and manage the tradeoff between privacy and convergence (Wei et al., 2020; Konečnỳ et al., 2016; Hamer et al., 2020; Lan et al., 2023). However, the utilization of differential privacy methods often results in substantial communication costs, particularly when applied to LLMs. In this article, we tackle this challenge through low-rank adaptation. The following section will provide more background of differential privacy.
3. Background of Differential Privacy
Deep neural networks. For simplicity, we start with a single-layer network, which takes an input vector $\bm{x}\in\mathbb{R}^{m}$ and transforms it into an output vector $\bm{\hat{y}}\in\mathbb{R}^{n}$ via a weight matrix $\bm{W}\in\mathbb{R}^{n\times m}$,
(1) $\bm{\hat{y}}=\bm{W}\cdot\bm{x}+\bm{b},$
where $\bm{b}\in\mathbb{R}^{n}$ is a bias vector. Let $\bm{\theta}$ denote the trainable parameters, i.e., $\bm{\theta}=\{\bm{W},\bm{b}\}$. A loss function $\mathcal{L}(\bm{\theta};\bm{x},\bm{y})$ is used to evaluate the difference between $\bm{\hat{y}}$ and $\bm{y}$, where $\bm{y}$ is the ground-truth label.
Assuming the training dataset $D$ has $N$ data pairs. During the training process, each iteration uses a batch to update the parameter $\bm{\theta}$, where each batch contains $B$ samples $\{(\bm{x}^{b},\bm{y}^{b})\}_{b=1}^{B}$. The parameter $\bm{\theta}$ is updated by the gradient descent method as follows
(2) $\bm{g}$ $=\frac{1}{B}\nabla_{\bm{\theta}}\sum\limits_{b=1}^{B}\mathcal{L}(\bm{\theta};\bm{x}^{b},\bm{y}^{b}),$
(3) $\bm{\theta}$ $=\bm{\theta}-\gamma\cdot\bm{g},$
where $\gamma$ is the learning rate. The training process executes for $T$ iterations.
Distributed training. We consider distributed training of a deep neural network with parameter $\bm{\theta}$. There are $K$ nodes participating in distributed training. Each node $k$ performs the gradient descent using its private dataset $D_{k}$, which contains $N_{k}$ samples. Thus, for node $k$, we have the sampling probability $q_{k}=B/N_{k}$. There are $N$ samples in total, where $N=\sum_{k=1}^{K}N_{k}$. In particular, one distributed training iteration consists of the following three:
  1. 1)
    A server broadcasts a global model $\bm{\theta}$ to all $K$ nodes.
  2. 2)
    Node $k$ initializes $\bm{\theta}_{k}$ by $\bm{\theta}$ and updates $\bm{\theta}_{k}$ by (3), using a batch of $B$ samples from $D_{k}$.
  3. 3)
    The server receives $\{\bm{\theta}_{1},...,\bm{\theta}_{K}\}$ from all $K$ nodes and updates $\bm{\theta}$ by
    (4) $\bm{\theta}=\sum\limits_{k=1}^{K}\rho_{k}\cdot\bm{\theta}_{k},$
where the $\rho_{k}$ is the weight factor, i.e., $\sum_{i=1}^{K}\rho_{k}=1$. The training process executes for $T$ iterations. The training process executes for $T$ iterations.
Differential privacy. However, the distributed training comes with a privacy issue (Carlini et al., 2021). Some attackers surveil the communication between the server and nodes and infer the information in dataset $D_{k}$. We need to provide a privacy guarantee for neural network training. Differential privacy considers the divergence of two distributions. There are multiple options to define the distance between distributions, and KL divergence is a popular choice. On the other hand, a $(\epsilon,\delta)$-differential private mechanism $\mathcal{M}$ is defined as
(5) $\text{Pr}[\mathcal{M}(d)=o]\leq e^{\epsilon}~{}\text{Pr}[\mathcal{M}(d^{\prime})=o]+\delta,$
where $d$ and $d^{\prime}$ are two adjacent inputs and $o$ stands for the output of mechanism $\mathcal{M}$. The divergence is bounded by $\epsilon$, and a relaxation of $\delta$ is allowed. Thus, we introduce randomness into the training process.
DP-SGD in (Abadi et al., 2016) proposes to add Gaussian noise into the gradient and provides a privacy guarantee for network training. For each iteration, the gradient $\bm{g}$ is replaced by:
(6) $\widetilde{\bm{g}}=\frac{\bm{g}}{\max(1,\frac{\bm{||\bm{g}||_{2}}}{C})}+\mathcal{N}(0,\sigma^{2}C^{2}\bm{I}),$
where $C$ is the clipping bound and $\mathcal{N}(0,\sigma^{2}C^{2}\bm{I})$ is the Gaussian distribution with mean 0 and standard deviation $\sigma^{2}C^{2}\bm{I}$. Under the Gaussian mechanism, each iteration is differentially private.
The training process can be seen as a composition of $T$ differentially private algorithms.
Theorem 1 ().
(Sequential composition (McSherry, 2009)) Let $\mathcal{M}_{k}$ each provides $(\epsilon_{k},\delta_{k})$-differential privacy, $k=1,2,...,K$. The sequential composition of $\mathcal{M}_{k}$ provides $(\epsilon_{k}^{{}^{\prime}},\delta_{k}^{{}^{\prime}})$-differential privacy, where the $\epsilon_{k}^{{}^{\prime}}=\sum_{k}\epsilon_{k}$ and $\delta_{k}^{{}^{\prime}}=\sum_{k}\delta_{k}$.
In Theorem 2, we can be shown that the training process essentially follows the differential privacy mechanism, and by incorporating moment counting, the noise has a scale of $\sigma$.
Theorem 2 ().
(Abadi et al., 2016) There exists constant $c_{1}$ and $c_{2}$ so that with the sampling probability $q=B/N$ over $T$ iterations, for any $\epsilon<c_{1}q^{2}T$, the network training process using DP-SGD method is $(\epsilon,\delta)$-differentially private for a given $\delta>0$ as long as
(7) $\sigma\geq c_{2}\frac{q\sqrt{T\log(1/\delta)}}{\epsilon}.$
Figure 1. Differential private low-rank adaptation scheme for LLMs using federated learning. Each node communicates with the server in updating $\bm{\theta}$.
4. Differentially Private Low-Rank Adaptation (DP-LoRA) of LLMs
Current advancements in large language models have led to an increased focus on open-source models by various research institutions, such as the Linux Foundation, Hugging Face, and AI4Finance Foundation. These organizations aim to foster large language models in vertical domains through collaborative efforts. For instance, banks are coming together to train financial models on transaction strategies and risk management and hospitals collaborate to develop models for diagnosing and treating medical conditions. These collaborations often involve fine-tuning existing large language models with private data to adapt them to specific tasks.
Each participating institution in such collaborations possesses sensitive, non-public data. Due to the constraints of privacy laws, this data cannot be shared among collaborating entities. As a result, institutions adopt the federated learning approach, in which each participant trains large language model locally and shares only the model weights in collaboration. However, studies have discovered that it is possible to extract training data from the weights of large language models. This data extraction attack means that the current method of sharing model weights in collaborations still poses a risk of private data leakage. Given these risks, there is an urgent need to establish a mechanism that provides privacy protection in collaborative model development.
We propose a differentially private low-rank adaptation (DP-LoRA) method of LLMs in federated learning. We adopt a Gaussian mechanism that adds noise to the model weights during the training process. Our method introduces a privacy guarantee for low-rank adaptation which can also reduce the communication overhead in federated learning. First, we discuss differential private low-rank adaptation in each node. We then address the communication scheme between each node and the server.
4.1. Problem Formulation
In this subsection, we formally describe the problem of differentially private distributed training for LLMs. In what follows, we first introduce differentially private distributed training and analyze the communication overhead. Then we describe our objective formally.
Differentially private distributed training. DP-SGD can also generalize to distributed training. Each node $k$ adopts (6) to calculate gradient $\widetilde{\bm{g}_{k}}$ and updates $\bm{\theta}_{k}$ by (3). Therefore, in node $k$, we can guarantee the ($\epsilon_{k},\delta_{k}$)-differential privacy for one iteration by adding Gaussian noise with mean 0 and standard deviation $\sigma\geq C\sqrt{2\ln(1.25/\delta_{k})}/\epsilon_{k})$.
Theorem 1 ().
(Parallel composition (McSherry, 2009)) Let $\mathcal{D}$ be a domain, let $D\in\mathbb{N}^{\mathcal{D}}$ be a data set, and let $K$ be a positive integer. For $k=1,2,...,K$, let $D_{k}$ be a subset of $\mathcal{D}$, let $\mathcal{M}_{k}$ be a mechanism that takes $D\cap D_{k}$ as input, and suppose $\mathcal{M}_{k}$ is $(\epsilon_{k},\delta_{k})$-differentially private. If $D_{i}\cap D_{j}=\emptyset$ whenever $i\neq j$, then the composition of the sequence $\mathcal{M}_{1},\ldots,\mathcal{M}_{k}$ is $(\max\{\epsilon_{i}\},\max\{\delta_{i}\})$-differentially private, where $i=1,...,k$.
As each $D_{k}$ came from a different distribution, the composition of $K$ differential private mechanism satisfies $(\max(\epsilon_{k}),\max(\delta_{k}))$-differential private. Therefore, one iteration of global model updates is a differentially private mechanism. The whole training process conducts $T$ iterations. Following Theorem 7, we can have the noise scale to guarantee the whole training process as a differential private mechanism, which can be represented as $\Omega(\frac{\max(q_{k})\sqrt{T\log(1/\delta)}}{\epsilon})$, where $q_{k}=B/N_{k}$.
Communication overhead. The communication overhead in federated learning involves several factors that contribute to the amount of data transmitted and the efficiency of this transmission. The communication overhead in federated learning is mainly related to five elements, which are model size, number of iterations, number of participants, and network conditions. The number of participants and update frequency can be adjusted easily and we assume the network condition is ideal without any noise. Thus, in this paper, we assume the communication overhead is mainly determined by the transferring model size. Unfortunately, LLMs often have a large number of parameters, which incurs significant communication overhead with federated learning.
Differentially private distributed training for LLMs. Given $K$ parties (e.g., hospitals or banks), where each party owns a private text dataset $D_{k}$ ($k=1,2,...,K$) for fine-tuning an LLM. We aim to develop a differentially private distributed training algorithm for the LLM, such that (1) the ($\epsilon_{k},\delta_{k}$)-differential privacy can be guaranteed, and (2) the communication overhead is minimized.
4.2. Proposed DP-LoRA Algorithm
Fig. 1 shows the architecture of our proposed DP-LoRA algorithm. Every node $k$ has a local dataset $D_{k}$. The dataset $D_{k}$ is not shared with the server or other nodes and has $N_{k}$ samples. We use $N$ to denote the total number of samples, where $N=\sum\limits_{k=1}^{K}N_{k}$. For the large language model, each node $k$ freeze the original model weight $\bm{W}$ and fine-tune the $\bm{\theta}_{k}$, where $\bm{\theta}_{k}$ is composed by low-rank decomposition matrix of each attention head. For each transformer layer $l$, the attention weight $\bm{W}^{l}$ is decomposed to $\bm{A}^{l}$ and $\bm{B}^{l}$. Therefore, $\bm{\theta}=\{\bm{A}^{1},\bm{B}^{1};\bm{A}^{2},\bm{B}^{2};...;\bm{A}^{L},\bm{B}^{L}\}$.
  • 1
    Broadcast: The server sends the global model parameters $\bm{\theta}$ to all participating nodes. This is the starting point for training in each round.
  • 2
    Local Training: Each node $k$ trains the model locally on its own dataset $D_{k}$. The freeze $\bm{W}$ indicates that the original model only does inference during local training and the weights are not updated by the gradient. We obtain a new $\bm{\theta}_{k}$ using gradient descent.
  • 3
    Add Noise: A Gaussian noise $n_{k}\sim\mathcal{N}(0,\sigma^{2})$ is add to the $\bm{\theta}_{k}$ such that $\widetilde{\bm{\theta}_{k}}=\bm{\theta}_{k}+n_{k}$.
  • 4
    Upload: Every node $k$ uploads $\widetilde{\bm{\theta}_{k}}$ to the server.
  • 5
    Aggregate: The server aggregates the $\widetilde{\bm{\theta}_{k}}$ from all nodes and takes the weighted average $\sum\limits_{k=1}^{K}\rho_{k}\widetilde{\bm{\theta}_{k}}$ to updates $\bm{\theta}$, where $\rho_{k}=N_{k}/N$.
We elaborate step $2$ and step $3$ by Alg. 1, which can be exemplified as DP-LoRA. Given a local dataset $D_{k}$, a noise scale $\sigma$, a clipping norm $C$, a number of iterations $T$, a learning rate $\gamma$, and a batch size $B$, we initialize local model parameters $\theta_{k}$. The general process can be described as
  1. (1)
    Select a random batch: Choose $B$ samples from the dataset $D_{k}$ randomly.
  2. (2)
    Compute gradients: Calculate the gradients of the loss function $\mathcal{L}$ with respect to the parameters $\theta_{k}$, denoted as $\bm{g}_{k}$.
  3. (3)
    Clip gradients: Perform gradient clipping by scaling down the gradients if their L2 norm is greater than $C$, $\bm{g}_{k}\leftarrow\bm{g}_{k}/\max(1,\frac{\|\bm{g}_{k}\|_{2}}{C})$.
  4. (4)
    Add noise: Introduce Gaussian noise to the clipped gradients to ensure differential privacy, $\bm{g}_{k}\leftarrow\bm{g}_{k}+\mathcal{N}(0,\sigma^{2}C^{2}\bm{I})$, $k=1,2,...,K$.
  5. (5)
    Gradient descent update: Update the local model parameters via gradient descent, $\widetilde{\theta_{k}}\leftarrow\theta_{k}-\gamma{\bm{g}_{k}}$, where $\bm{g}_{k}$ is the noised and clipped gradient.
4.3. Guarantee of Locally Differential Privacy
We assume that there are $K$ nodes in total. Each node $k$ has a private dataset $D_{k},k=1,2,...,K$. Each node holds an initial weight $\bm{W}$ which is freezed and only used for making inferences. All nodes hold the same $\bm{W}$. The node $k$ uses the dataset $D_{k}$ in fine-tuning the low-rank adaptation modules with parameter $\bm{\theta}_{k}$.
We introduce a differentially private low-rank adaptation (DP-LoRA) as a local training algorithm. We add noise to the gradient in each iteration. For $T$ training iterations, the proposed algorithm can satisfy $(\epsilon,\delta)$-differential privacy. Our algorithm is presented as follows.
Algorithm 1 DP-LoRA Algorithm
1:  Input: number of nodes $K$, dataset $\{D_{k}\}_{k=1}^{K}$, noise scale $\sigma$, clipping norm $C$, iterations $T$, learning rate $\gamma$, batch size $B$, network parameter $\bm{\theta}$.
2:  Initialize $\bm{\theta}$
3:  for $t=0$ to $T$ do
4:     Send $\bm{\theta}$ to all $K$ nodes independently
5:     $\bm{\theta}_{k}\leftarrow$ Nodeupdate($D_{k}$, C, B, $\sigma$, $\gamma$),   $k=1,2,...,K$
6:     $\bm{\theta}\leftarrow\sum\limits_{k=1}^{K}\rho_{k}\bm{\theta}_{k}$
7:  end for
8:  Output: $\bm{\theta}$
9:
10:  function Nodeupdate($D_{k}$, C, B, $\sigma$, $\gamma$)
11:  sample $(\bm{x},\bm{y})$ from $D_{k}$ with batch size $B$
12:  $\bm{g}_{k}\leftarrow\nabla_{\bm{\theta}_{k}}\mathcal{L}(\bm{\theta}_{k};\bm{x},\bm{y})$
13:  $\bm{g}_{k}\leftarrow\bm{g}_{k}/\max(1,\frac{\|\bm{g}_{\bm{\theta}_{k}}\|_{2}}{C})$
14:  ${\bm{g}}_{k}\leftarrow{\bm{g}}_{k}+\mathcal{N}(0,\sigma^{2}C^{2}\bm{I})$
15:  $\bm{\theta}_{k}\leftarrow\bm{\theta}_{k}-\gamma\bm{g}_{k}$
16:  Return $\bm{\theta}_{k}$
17:  end function
A local model is composed by $L$ weight matrix $\bm{W}^{l}\in\mathbb{R}^{n\times n},l=1,2,...,L$. Here, we consider the linear layers, and the input of output shares the same dimension $n$. These weight matrices have a low “intrinsic dimension" (Aghajanyan et al., 2020). We decompose the matrix $\bm{W}^{l}$ into $\bm{A}^{l}\in\mathbb{R}^{n\times r}$ and $\bm{B}^{l}\in\mathbb{R}^{r\times n}$, follow by low-rank adaptation method (Hu et al., 2021). Note that the rank $r\ll n$. Thus, we have $\bm{\theta}=\{\bm{A}^{1},\bm{B}^{1};\bm{A}^{2},\bm{B}^{2};...;\bm{A}^{L},\bm{B}^{L}\}$. We follow the setting in (Abadi et al., 2016) to sample $B$ data from the dataset to provide an unbiased estimation.
Thirdly, we add Gaussian noise in the gradient. The probability density function for Gaussian noise is $\mathcal{N}(0,\sigma^{2}C^{2}\bm{I})$, where $\bm{I}$ is an identity matrix and has the same shape with $\bm{\theta}_{k}$. The clipping bound $C$ is used to ensure the gradient norm $||\bm{g}_{k}||<C$. The noise scale $\sigma$ provides the privacy guarantee for each iteration, which should follow $\sigma\geq\sqrt{2\ln{(1.25/\delta)}}/\epsilon$. We update the add noise $\mathcal{N}(0,\sigma^{2}C^{2}\bm{I})$ into matrix $\bm{A}$ and $\bm{B}$, respectively.
Equation (6) guarantees that one training iteration is $(\epsilon,\delta)$-differential privacy. Alg. 1 composes $T$ iterations. As the privacy amplification theorem (Hu et al., 2021; Beimel et al., 2014) and strong composition theorem (Dwork et al., 2014), the composition of $T$ differential private mechanism is still differential private. The algorithm 1 is $(\epsilon,\delta)$-differential private.
4.4. Reducing Communication Overhead in Federate Learning
The weight $\bm{\theta}$ is composed by $\bm{A}$ and $\bm{B}$. Federated learning involves collecting local model weights for $T$ iterations and redistributing the aggregated weights $\bm{\theta}$ back to each participating node. It is assumed that all nodes utilize Algorithm 1 for training and share identical hyper-parameters. The system comprises $K$ nodes in total. Each node possesses its private dataset $D_{k}$, where $k=1,2,\ldots,K$, and maintains local model weights $\bm{\theta}_{k}$, for $k=1,2,\ldots,K$. In Algorithm 1, each node $k$ refines its $\bm{\theta}_{k}$ using the corresponding dataset $D_{k}$. A central server maintains the global model weights $\bm{\theta}$.
We assess the reduction in communication overhead achieved through the low-rank adaptation method. Assuming that the federated learning process spans $T$ rounds in a noiseless and ideal network environment, the communication overhead can be quantified using specific parameters. The proposed method provides the reduction in two steps. First, we freeze most of the model weight. We fine-tune and transfer the weight in self-attention layers. For example, a Llama-7B is composed of $32$ transformer layers and each transformer consists $32$ attention heads, which have $4096$ dimensions. Thus, each self-attention block contains $4096\times 4096\times 3=50,331,648$ parameters, and $32$ transformers contain $1,610,612,736$ parameters in total. Therefore, we reduce the training parameter from $6,738,411,520$ to $1,610,612,736$. Second, we decomposed the self-attention parameters by low-rank adaptation methods. We use a $\bm{W}^{l}$ to denote the parameters of one attention head. Each $\bm{W}^{l}\in\mathbb{R}^{n\times n}$ is replace by a $\bm{A}^{l}\in\mathbb{R}^{n\times r}$ and $\bm{B}^{l}\in\mathbb{R}^{n\times r}$. Thus, for one attention head with $4096$ dimension, we reduce the parameter from $16,777,216$ to $2,097,152$ if we choose $r$ as $256$.
In general, the communication overhead is determined by the number of nodes $K$, the training epoch $T$, and the transmission parameters $\bm{\theta}$. The total communication overhead is given by $T\times K\times L\times r\times 2n$. Take training Llama-7B with 5 nodes in 50 iterations for example, the original communication overhead is $2\times 5\times 50\times 6.7B$ and the communication overhead for DP-lora is $2\times 5\times 50\times 4096\times 256$.
5. Performance Evaluation
5.1. Tasks and Datasets
To mimic the specific implementation scenario, we chose three datasets to test our algorithm. To mimic the protection of highly sensitive data, we choose the financial and medical data which always contains sensitive information. We also test our algorithm on a normal dataset to provide a comparison with other methods.
SlimPajama: open-source dataset for training large language models. Slimpajama is created from RedPajama (Computer, 2023) by filtering out low-quality data and duplicates. SlimPajama contains 627B tokens. The data in Slimpajama mainly comes from a wide array of online sources. (1). Arxiv: ArXiv contains preprints in fields like physics, mathematics, and computer science, offering early access to scholarly research articles before peer review. (2). Github: GitHub contains tools for code hosting, collaboration, version control, and project management for software development. (3). Wikipedia: Wikipedia contains a vast range of information, including articles on topics like history, science, geography, technology, culture, biographies, and much more, all written and continually updated by volunteer contributors from around the world.
Medical dataset: dataset contains 0.26 million English consultations between patients and doctors. The total number of utterances is 0.51 million. Each consultation consists of two parts: (1) a description of the patient’s medical conditions; (2) a conversation between the patient and the doctor. The data is crawled from "iclinic.com3" and "healthcaremagic.com", which are two online platforms of healthcare services, including symptom self-checker, video consultation, online chat with doctors, etc. The consultations cover 51 categories of communities including diabetes, elderly problems, pain management, etc., and 96 specialties including andrology, cardiology, nephrology, pharmacology, etc. The consultations were conducted from 2008 to 2020.
Finance dataset: We collect extensive financial data from various sources. (1). Financial news: Websites such as Reuters, CNBC, and Yahoo Finance, among others, are rich sources of financial news and market updates. (2). Social media: Platforms such as Twitter, Facebook, Reddit, Weibo, and others, offer a wealth of information in terms of public sentiment. (3). Filings: Websites of financial regulatory authorities, such as the SEC in the United States, offer access to company filings. (4). Trends: Websites like Seeking Alpha, Google Trends, and other finance-focused blogs and forums provide access to analysts’ opinions, and market predictions. (5). Academic datasets: Research-based datasets that offer curated and verified information for sophisticated financial analysis. The dataset comprises 9,540 samples for training, each annotated with one of three labels: Bearish, Bullish, or Neutral.
Table 1. Statistics of the training dataset
(a) Statistics of the Medical Dialogue dataset.
Metric Value
# dialogues 257,332
# utterances 514,664
# tokens 44,527,872
Avg. # of utterances in a dialogue 2
Max # of utterances in a dialogue 2
Min # of utterances in a dialogue 2
Avg. # of tokens in an utterance 86.5
Max # of tokens in an utterance 3,672
Min # tokens in an utterance 1
(b) Statistics of the Slimpajama.
Data source SlimPajama
Commoncrawl 52.2%
C4 26.7%
GitHub 5.2%
Books 4.2%
ArXiv 4.6%
Wikipedia 3.8%
StackExchange 3.3%
5.2. Experiments Setup
Evaluation Task: We assess our methods using test sets across different fields, aligned with the training data.
For general tasks, we use three datasets: BoolQ, a question-answering dataset with 16K examples, each containing a passage, a related yes/no question, and an answer. PIQA, focused on physical commonsense reasoning, comprising 21K questions with two solution options, where the model selects the most appropriate. WinoGrande, a commonsense reasoning dataset with 44K sentences, each with a blank to fill using one of two provided options.
In the financial domain, we use FPB, with 4.8K financial news sentences categorized by sentiment (positive, negative, neutral). FiQA SA, comprising 17K microblog headlines and financial news sentences, is classified by sentiment. TFNS, a collection of 11.9K finance-related tweets, is used for sentiment classification.
For medical tasks, our datasets include MedQuAD, which offers 47,457 medical question-answer pairs from 12 NIH websites, covering 37 question types. LiveQA Test, featuring questions across 26 types and five categories, each with subquestions, a focus, and validated reference answers from trusted resources. MEDIQA-Ans, containing 156 medical questions.
Models: To provide a comprehensive understanding of our algorithm, we test our algorithm based on different language models. (1) GPT-2(Radford et al., 2019): We chose the basic version of GPT-2 in our experiments. It has 1.5 billion parameters and it is composed of 12 layers and each layer has 768 dimensions. The model is released on (Radford et al., 2019). (2) Bert (Devlin et al., 2018): We chose the Bert base version to conduct our experiments. The Bert base uses 12 layers and 12 attention heads and has a total number of Parameters of 110M. The pre-trained weight can be found in the Link from (Devlin et al., 2018). (3) ChatGLM-6b (Zeng et al., 2022): ChatGLM2-6B has 6.2 billion parameters. It contains 28 layers with dimensions of 4096 and heads of 32. (4) Llama2-7B (Touvron et al., 2023): Llama2-7B contains 6.7B parameters. It contains 32 layers with dimensions of $4096$ and attheads of $32$. The initial Llama-2 weights are downloaded from the given link (Touvron et al., 2023).
Experiments settings: For federated learning, We split the dataset evenly and stored it locally among all $K$ nodes. We adopt an average for aggregating local model weights and we set all $\rho_{k}$ in Equation 4 to $\frac{1}{K}$. We assume that the communication channels are noiseless and ideal. Each node $k$ uploads its $\bm{\theta}_{k}$ to the server in every iteration. We choose $K$ as $5$ and we set the maximum iteration $T$ as 50. For the training, we follow the settings in 2.
Table 2. hyper-parameter settings for training
hyper-parameter Value
batch size $B$ $8$
noise scale $\sigma$ $2$
learning rate $\gamma$ $5e^{-4}$
clipping bound $C$ $10$
5.3. Results on Medical Tasks
In our experiments, we evaluated the performance of several models in the context of differential privacy for medical tasks. We investigated the impact of varying the parameters $\epsilon$ and $\delta$ for the language model fine-tuning on the medical datasets. In all experiments, the compression rank was fixed at $512$ and the $\delta$ is set to $51e-$ 5 when we varying the $\epsilon$ and the $\epsilon$ is settd to $2$.
One notable trend is the general decrease in performance across all models with stricter privacy settings (lower $\epsilon$ and higher $\delta$ values). For instance, GPT-2’s performance on MedQuAD drops from 69.2 at the original setting to 55.3 and 49.1 when $\epsilon$ is reduced to 2 and $\delta$ to $051e-$ 05, respectively (Tables 3 and 7). This trend indicates a trade-off between privacy and utilities.
Table 3. The performance of the global model under different $\epsilon$ for GPT-2 in general and Medical tasks.
$\epsilon$ BoolQ PIQA WinoGrande MedQuAD LiveQA MEDIQA-Ans
origin 61.9 53.8 48.3 69.2 39.5 55.1
2 51.5 46.3 46.1 55.3 27.0 49.1
4 49.0 45.4 46.7 57.3 37.1 47.4
6 58.4 41.5 45.6 63.5 38.9 52.9
8 52.9 45.1 42.7 66.6 39.1 47.9
10 58.9 53.1 44.7 64.1 35.9 51.8
Table 4. The performance of the global model under different $\epsilon$ for Bert in general and Medical tasks.
$\epsilon$ BoolQ PIQA WinoGrande MedQuAD LiveQA MEDIQA-Ans
origin 45.8 51.3 41.2 69.1 56.4 73.3
2 36.2 41.0 29.8 53.2 46.8 66.5
4 33.1 37.2 38.0 63.6 46.3 61.0
6 40.1 40.2 33.1 68.8 47.6 60.9
8 40.1 42.2 38.9 60.3 51.5 70.2
10 42.2 45.9 37.9 68.9 51.6 69.2
Table 5. The performance of the global model under different $\epsilon$ for ChatGLM-6B in general and Medical tasks.
$\epsilon$ BoolQ PIQA WinoGrande MedQuAD LiveQA MEDIQA-Ans
origin 44.2 51.2 50.1 66.6 71.9 69.3
2 34.0 47.9 34.9 66.4 67.3 57.8
4 32.9 45.4 46.0 57.1 57.5 62.2
6 37.8 42.2 47.5 60.6 58.9 65.8
8 37.4 47.2 48.7 59.8 69.5 66.9
10 38.8 47.9 43.8 65.9 69.8 67.7
Table 6. The performance of the global model under different $\epsilon$ for Llama-7B in general and Medical tasks.
$\epsilon$ BoolQ PIQA WinoGrande MedQuAD LiveQA MEDIQA-Ans
origin 71.3 66.2 65.5 73.1 69.4 65.5
2 54.9 57.1 53.4 66.5 55.9 52.8
4 60.6 59.5 55.5 66.6 67.3 53.2
6 60.8 64.9 51.7 67.3 68.1 65.3
8 66.3 64.1 60.4 66.2 60.5 62.1
10 66.6 62.2 65.2 69.7 66.1 58.6
In contrast, some models like Llama-7B show a more resilient performance under varying $\epsilon$ values. For example, its performance on LiveQA only marginally decreases from 69.4 to 66.1 when $\epsilon$ is increased from the original to 10 (Table 6). This suggests that certain models might be better suited for privacy-sensitive applications.
Additionally, the impact of changing $\delta$ values appears to be more model-specific. Bert’s performance on MEDIQA-Ans decreases significantly from 73.3 in the original setting to 57.4 when $\delta$ is reduced to $061e-$ 06 (Table 8), highlighting a potentially higher sensitivity to $\delta$ adjustments.
Table 7. The performance of the global model under different $\delta$ for GPT-2 in general and Medical tasks.
$\delta$ BoolQ PIQA WinoGrande MedQuAD LiveQA MEDIQA-Ans
origin 61.9 53.8 48.3 69.2 39.5 55.1
1e-02 60.2 50.7 46.6 69.1 32.5 51.1
1e-03 54.7 49.5 38.7 59.8 36.1 46.1
1e-04 61.6 47.6 35.0 68.5 30.9 50.3
1e-05 51.5 46.3 46.1 55.3 27.0 49.1
1e-06 53.4 42.9 40.6 69.1 30.5 41.4
Table 8. The performance of the global model under different $\delta$ for Bert in general and Medical tasks.
$\delta$ BoolQ PIQA WinoGrande MedQuAD LiveQA MEDIQA-Ans
origin 45.8 51.3 41.2 69.1 56.4 73.3
1e-02 37.7 44.3 37.8 65.3 50.9 73.3
1e-03 37.3 44.1 38.6 65.4 48.2 67.5
1e-04 35.1 40.1 33.6 64.3 52.5 59.1
1e-05 36.2 41.0 29.8 53.2 46.8 66.5
1e-06 31.3 43.2 36.4 62.9 40.6 57.4
Table 9. The performance of the global model under different $\delta$ for ChatGLM-6B in general and Medical tasks. The left column denotes the maximum value of $\delta$.
$\delta$ BoolQ PIQA WinoGrande MedQuAD LiveQA MEDIQA-Ans
origin 44.2 51.2 50.1 66.6 71.9 69.3
1e-02 37.7 44.3 37.8 65.3 50.9 73.3
1e-03 43.3 44.1 38.6 65.4 48.2 67.5
1e-04 39.1 40.1 33.6 64.3 55.5 59.1
1e-05 34.0 47.9 34.9 66.4 67.3 57.8
1e-06 31.3 43.2 36.4 62.9 40.6 57.4
Table 10. The performance of the global model under different $\delta$ for Llama-7B in general and Medical tasks. The left column denotes the maximum value of $\delta$.
$\delta$ BoolQ PIQA WinoGrande MedQuAD LiveQA MEDIQA-Ans
origin 71.3 66.2 65.5 73.1 69.4 65.5
1e-02 68.7 60.3 63.7 70.6 68.7 60.2
1e-03 67.1 62.6 65.0 64.1 63.8 62.1
1e-04 59.9 57.9 56.6 70.3 60.7 58.0
1e-05 54.9 57.1 53.4 66.5 55.9 52.8
1e-06 49.7 48.4 59.1 59.9 49.3 44.3
We can observe a significant trade-off between privacy and utility in medical tasks using models like GPT-2, Bert, ChatGLM-6B, and Llama-7B. Our data clearly shows that stricter privacy settings (lower $\epsilon$ and $\delta$ values) generally correlate with reduced performance across tasks.
5.4. Results on Financial Tasks
In this section, we compare the financial tasks performance of the global model which is fine-tuned on the financial datasets. In our experiments, We assume that there are $5$ clients in participating the whole training process. We use the above-mentioned models as our initial models. All the local models and global models follow the same size and use the original model weights in the initial. We trained each local model under the same set of parameters, such as learning rate, decay strategy, and training epochs. We set the compression rank $r$ to 512 and $\epsilon$ value as 2 when we vary the $\delta$ and $\delta$ as $51e-$ 5 when we vary the $\epsilon$. The origin row means that we train our model without differential private adaptations.
Table 11. The performance of the global model under different $\delta$ for GPT-2 in general and financial tasks. The left column denotes the maximum value of $\delta$.
$\delta$ BoolQ PIQA WinoGrande FPB FiQA SA TFNS
origin 41.1 39.9 37.5 59.8 73.8 66,0
1e-02 36.4 39.6 35.7 57.1 71.4 57.0
1e-03 34.7 35.4 29.4 51.1 63.3 52.2
1e-04 33.1 28.6 27.8 42.1 57.5 45.1
1e-05 25.5 20.5 22.2 39.1 42.9 34.3
1e-06 23.2 19.0 20.3 28.1 52.2 34.0
Table 12. The performance of the global model under different $\delta$ for Bert in general and financial tasks. The left column denotes the maximum value of $\delta$.
$\delta$ BoolQ PIQA WinoGrande FPB FiQA SA TFNS
origin 45.3 47.5 51.3 65.5 60.1 57.3
1e-02 41.8 41.7 46.7 59.6 59.1 55.6
1e-03 37.9 39.7 44.9 50.6 55.6 50.9
1e-04 37.7 36.8 39.2 48.4 51.4 47.9
1e-05 21.7 29.9 38.8 43.7 47.5 44.1
1e-06 20.3 28.7 30.2 37.3 45.3 40.3
Table 13. The performance of the global model under different $\delta$ for ChatGLM-6B in general and financial tasks. The left column denotes the maximum value of $\delta$.
$\delta$ BoolQ PIQA WinoGrande FPB FiQA SA TFNS
origin 61.5 69.3 63.8 55.5 74.3 70.1
1e-02 55.5 68.7 61.1 50.7 69.4 62.2
1e-03 53.6 65.3 54.9 49.1 64.5 56.2
1e-04 50.7 62.3 51.7 47.4 58.9 47.6
1e-05 44.9 58.3 48.6 29.9 48.1 42.5
1e-06 38.5 55.2 47.4 30.6 51.5 38.0
Table 14. The performance of the global model under different $\delta$ for Llama-7B in general and financial tasks. The left column denotes the maximum value of $\delta$.
$\delta$ BoolQ PIQA WinoGrande FPB FiQA SA TFNS
origin 76.3 79.7 70.0 59.1 80.0 69.6
1e-02 88.1 80.6 69.3 80.3 81.9 67.3
1e-03 81.8 78.2 57.6 64.1 75.0 65.4
1e-04 75.4 69.3 59.1 49.1 76.1 70.3
1e-05 59.3 71.5 61.6 38.5 48.3 43.2
1e-06 69.5 56.6 63.5 35.2 51.5 28.9
Table 15. The performance of the global model under different $\epsilon$ for GPT-2 in general and financial tasks. The left column denotes the maximum value of $\epsilon$.
$\epsilon$ BoolQ PIQA WinoGrande FPB FiQA SA TFNS
origin 41.1 39.9 37.5 59.8 73.8 66
2 25.5 20.5 22.2 39.1 42.9 34.3
4 26.8 19.5 29.3 40.8 53.9 35.2
6 37.3 26.5 28.2 49.0 55.5 49.6
8 33.9 27.5 31.1 53.2 56.9 54.7
10 39.5 31.2 31.2 53.9 57.1 54.8
Table 16. The performance of the global model under different $\epsilon$ for Bert in general and financial tasks. The left column denotes the maximum value of $\epsilon$.
$\epsilon$ BoolQ PIQA WinoGrande FPB FiQA SA TFNS
origin 45.3 47.5 51.3 65.5 60.1 57.3
2 21.7 29.9 38.8 43.7 47.5 44.1
4 24.5 33.5 39.9 45.2 47.3 44.5
6 27.5 34.6 45.0 49.9 53.8 45.7
8 31.0 36.2 46.7 50.4 54.9 51.9
10 29.1 37.6 47.5 52.3 59.9 49.2
Table 17. The performance of the global model under different $\epsilon$ for ChatGLM-6B in general and financial tasks.. The left column denotes the maximum value of $\epsilon$.
$\epsilon$ BoolQ PIQA WinoGrande FPB FiQA SA TFNS
origin 61.5 69.3 63.8 55.5 74.3 70.1
2 44.9 58.3 48.6 29.9 48.1 42.5
4 43.2 58.3 48.3 40.2 57.9 54.8
6 48.4 59.1 51.4 43.5 62.2 56.3
8 52.5 60.9 58.0 45.6 68.1 63.2
10 55.3 67.0 62.2 53.2 72.6 64.9
Table 18. The performance of the global model under different $\epsilon$ for Llama-7B in general and financial tasks. The left column denotes the maximum value of $\epsilon$.
$\epsilon$ BoolQ PIQA WinoGrande FPB FiQA SA TFNS
origin 76.3 79.7 70.0 59.1 80.0 69.6
2 59.6 71.5 61.6 38.5 48.3 43.2
4 63.4 69.3 61.9 43.1 65.0 65.4
6 71.9 70.3 59.1 49.1 76.1 65.3
8 76.4 80.4 69.9 68.5 73.1 77.2
10 87.9 80.1 71.1 69.9 81.5 82.9
The experimental data consistently demonstrate a clear trade-off between the privacy level (as adjusted by $\epsilon$ and $\delta$ values) and the model performance across various tasks. While all models show a decrease in performance with stricter privacy settings, the extent of this decrease varies, suggesting differences in how each model adapts to privacy constraints.
One notable trend is the general decrease in performance across all models with stricter privacy settings (lower $\epsilon$ and $\delta$ values). For instance, GPT-2’s performance on MedQuAD drops from 69.2 at the original setting to 55.3 and 49.1 when $\epsilon$ is reduced to 2 and $\delta$ to $051e-$ 05, respectively (Tables 3 and 7). This trend indicates a trade-off between privacy and effectiveness.
In contrast, some models like Llama-7B show a more resilient performance under varying $\epsilon$ values. For example, its performance on LiveQA only marginally decreases from 69.4 to 66.1 when $\epsilon$ is increased from the original to 10 (Table 6). This suggests that certain models might be better suited for privacy-sensitive applications.
Additionally, the impact of changing $\delta$ values appears to be more model-specific. Bert’s performance on MEDIQA-Ans decreases significantly from 73.3 in the original setting to 57.4 when $\delta$ is reduced to $061e-$ 06 (Table 8), highlighting a potentially higher sensitivity to $\delta$ adjustments.
Table 19. We show the communication overhead of GPT-2 in federated learning. We use the model parameter amount to denote the communication efficiency and compression ratio as an improvement indicator for different sets of rank $r$. The origin row means the results without our algorithm. For all the experiments, we train our local model under privacy parameter $\epsilon=8$ and $\delta=10e^{-5}$. We also include the model performance on general tasks.
rank $r$ BoolQ PIQA WinoGrande Communication overhead Reduction ratio
origin 61.9 56.9 48.9 127.4M -
512 39.4 41.7 45.5 101.3M 79.5%
256 38.2 44.3 43.2 82.1M 64.4%
128 32.1 50.3 45.9 69.5M 54.6%
64 32.4 49.1 39.9 58.6M 46.0%
32 33.5 46.3 39.4 58.0M 45.5%
Table 20. We show the communication overhead of Bert in federated learning. We use the model parameter amount to denote the communication efficiency and compression ratio as an improvement indicator for different sets of rank $r$. The origin row means the results without our algorithm. For all the experiments, we train our local model under privacy parameter $\epsilon=8$ and $\delta=10e^{-5}$. We also include the model performance on general tasks.
rank $r$ BoolQ PIQA WinoGrande Communication overhead Reduction ratio
origin 49.8 55.5 60.2 110.8M -
512 35.5 44.7 41.3 83.7M 75.5%
256 34.2 47.3 40.3 69.9M 63.1%
128 35.1 46.2 43.8 54.5M 49.2%
64 33.2 39.1 38.9 48.1M 43.4%
32 33.7 37.0 37.8 46.6M 42.0%
Table 21. We show the communication overhead of Llama-7B in federated learning. We use the model parameter amount to denote the communication efficiency and compression ratio as an improvement indicator for different sets of rank $r$. The origin row means the results without our algorithm. For all the experiments, we train our local model under privacy parameter $\epsilon=8$ and $\delta=10e^{-5}$. We also include the model performance on general tasks.
rank $r$ BoolQ PIQA WinoGrande Communication overhead Reduction ratio
origin 76.3 79.7 70.0 6.7B -
1024 78.8 83.1 71.1 2.43B 36.27%
512 76.4 80.4 69.9 1.35B 20.15%
256 70.6 74.1 68.3 0.93B 13.88%
128 50.3 78.5 69.7 0.65B 9.70%
64 39.1 68.5 67.8 0.49B 7.31%
Table 22. We show the communication overhead of ChatGLM-6B in federated learning. We use the model parameter amount to denote the communication efficiency and compression ratio as an improvement indicator for different sets of rank $r$. The origin row means the results without our algorithm. For all the experiments, we train our local model under privacy parameter $\epsilon=8$ and $\delta=10e^{-5}$. We also include the model performance on general tasks.
rank $r$ BoolQ PIQA WinoGrande Communication overhead Reduction ratio
origin 69.5 76.3 71.3 6.2B -
1024 57.3 65.3 55.5 2.94B 47.4%
512 55.7 59.7 56.5 1.76B 28.4%
256 47.6 55.4 48.8 0.83B 13.4%
128 51.3 56.7 44.1 0.52B 9.3%
64 50.1 49.9 46.3 0.47B 7.6%
5.5. Reduction in Communication Overhead
We test the performance of the global model and communication overhead in federated learning. We choose the rank $r$ in a decreasing manner. We use model parameter amount to indicate communication overhead. The local models are trained with Alg. 1. We set the privacy parameters $\epsilon=8$ and $\delta=10e-5$.
Across all models, we observed a consistent trend: as the rank $r$ decreased, the number of parameters and consequently the communication overhead also decreased. This trend indicates a direct correlation between rank reduction and communication efficiency. However, there is a trade-off observed in model performance on general tasks for the differential privacy training.
GPT-2 (Table 19) showed a performance drop in BoolQ from 61.9 to 33.5, PIQA from 56.9 to 46.3, and WinoGrande from 48.9 to 39.4 as the rank reduced from origin to 32, alongside an increase in compression ratio from 45.5 to 79.5. BERT (Table 20) exhibited a similar performance decline, for example, in WinoGrande from 60.2 to 37.8, with a compression improvement from 42.0 to 75.5. Llama-7B (Table 21), a larger model, maintained a relatively stable performance (e.g., BoolQ at 76.3 to 39.1, WinoGrande 70.0 to 67.8 ) despite significant parameter reduction (compression ratio from 7.31 to 36.27). In contrast, ChatGLM-6B (Table 22) showed a more pronounced performance degradation (BoolQ dropping from 69.5 to 50.1, WinoGrande from 71.3 to 46.3) with a notable decrease in parameters (compression ratio moving from 7.6 to 47.4). These results highlight the variability in how different models respond to parameter reduction and compression, underscoring the challenge of balancing communication efficiency with performance in federated learning environments.
6. Conclusion
In this paper, we have introduced DP-LoRA, an innovative approach to preserve data privacy in the low-rank adaptation of LLMs. DP-LoRA ensures differential privacy while significantly reducing communication overhead. By introducing noise into gradients to achieve ($\epsilon$, $\delta$)-differential privacy, we enhance communication efficiency. Empirical validation through experiments demonstrates the efficacy of our approach, rendering LLMs more secure for domain-specific fine-tuning. Looking ahead, our future research endeavors will explore refining the error bounds for differential privacy, particularly in the context of fine-tuning LLMs where the existing bounds may be less stringent. Additionally, we aim to broaden our experimental scope to encompass a wider array of domain-specific datasets, further validating the applicability and robustness of DP-LoRA.